OpenStudy (anonymous):

Solve each equation for 0 degrees < x < 360 degrees (round to nearest tenth of degree). a) sin x = -1/3 b) cot x = -√3/3

OpenStudy (michele_laino):

please for your first equation, you have to use a scientific calculator, and keep in mind that there are two distinct solutions. For your second equation, please keep in minde that: \[\cos(120°)=-\frac{ 1 }{ 2 }, \sin(120)=\frac{ \sqrt{3} }{ 2 }\]